# Energy-momentum tensor in the 2D Ising CFT in full modular space

### v1

@@ 1 Introduction
To expand the non-perturbative understanding of quantum field theory, the Quantum Finite Elements (QFE) project [Brower:2016moq] aims to enable lattice calculation on curved spacetime. Such an extension enables us to, for example, use radial quantization for conformal field theories (CFTs) in the spacetime with dimensions $D>2$ [Fubini:1972mf, Cardy:1985lth, Brower:2012vg]. As a step towards its rigorous formulation, this work considers the energy-momentum (EM) tensor in the 2D CFT on the torus with generic modulus $\tau$, which is a key physical object that generates diffeomorphism. We derive a lattice expression of the EM tensor in both fermionic and spin variables. The expression is confirmed in direct Monte Carlo calculation by checking the conformal Ward identities [Eguchi:1986sb, DiFrancesco:1987ez]. Although we are not the first people who address the EM tensor in the two-dimensional Ising spin model [Kadanoff:1970kz], we derive a complete expression including normalization that works under an arbitrary affine transformation. To our knowledge, our work is the first work that verifies the analytic formulas for the Ising CFT involving the EM tensor numerically with the spin variables. The details of the content will be given in a subsequent paper [NM2025].
@@ 3 Partition functions
As is well known, the 2D Ising model has a fermionic description with the Majorana fermion [PhysRev.65.117, RevModPhys.36.856, Itzykson:1982ed]. For later discussion, we consider the Wilson-Majorana fermion action [Wolff:2020oky, Brower:2022cwv] on the hexagonal lattice in a generalized setup with arbitrary couplings:
@@
The partition functions (4), (5) and (6) can be related to each other by loop expansions [10.1063/1.524404, Wolff:2020oky, Brower:2022cwv] even under various boundary conditions on the torus. For the discussion below, it suffices to quote the following identities:
@@ 4 Staggered lattice structure in the affine-transformed hexagonal lattice
In Ref. [Brower:2022cwv], it was shown that the continuum limit can be taken under nontrivial affine transformation with the couplings:
@@ 5 Lattice operators
We comment that both $\varepsilon(x)$ and $T_{k}(x)$ acquire divergent parts through loop diagrams (the divergent part of $\varepsilon(x)$ comes from the Wilson term). The divergent parts need to be subtracted when we consider disconnected components. This corresponds to taking the normal ordering. For the Ising CFT, its evaluation can be performed with the fermion system without statistical error [NM2025]. In the numerical analysis below, all the lattice operators are regularized in this way.
@@ 6 Confirming the CFT formulas with Monte Carlo calculations
In this section, we calculate various expectation values that involve the EM tensor with the spin variable by performing Monte Carlo calculation. Ensembles are generated by combining the Wolff algorithm [Wolff:2020oky] with the local heat-bath algorithm. The details of the calculation will be described in Ref. [NM2025]. We comment that that the signal of the EM tensor is noisy as the contributions from 1 (the diverging part), $\varepsilon(x)$, and $\partial\varepsilon(x)$ (the first descendant of $\varepsilon(x)$) need to be subtracted to single out the EM tensor components. The exact values are evaluated with the analytic expressions known in the Ising CFT [Eguchi:1986sb, DiFrancesco:1987ez]. The modulus is set to a nontrivial value $\tau=1.2e^{4i\pi/9}$ below.
@@
We next consider the three-point function with the spin operators. This is the least noisy correlator involving the EM tensor. Recall that the conformal Ward identity for the primary fields $\phi_{i}(z,\bar{z})$ with the conformal weights $(h_{i},\tilde{h}_{i})$ is [Eguchi:1986sb]:
@@ Acknowledgments
The authors thank Hidenori Fukaya, Okuto Morikawa, and Yusuke Namekawa for valuable discussions in LATTICE 2024. This work is partially supported by the Scientific Discovery through Advanced Computing (SciDAC) program, “Multiscale acceleration: Powering future discoveries in High Energy Physics” under FOA LAB-2580 funded by the U.S. DOE, Office of Science, and DOE under Award DE-SC0015845. The computation was performed on the Shared Computing Cluster (SCC) which is administered by Boston University’s Research Computing Services [SCC].
[…]

### v2

@@ 1 Introduction
To expand the non-perturbative understanding of quantum field theory, the Quantum Finite Elements (QFE) project [1] aims to enable lattice calculation on curved spacetime. Such an extension enables us to, for example, use radial quantization for conformal field theories (CFTs) in the spacetime with dimensions $D>2$ [2, 3, 4]. As a step towards its rigorous formulation, this work considers the energy-momentum (EM) tensor in the 2D CFT on the torus with generic modulus $\tau$, which is a key physical object that generates diffeomorphism. We derive a lattice expression of the EM tensor in both fermionic and spin variables. The expression is confirmed in direct Monte Carlo calculation by checking the conformal Ward identities [5, 6]. Although we are not the first people who address the EM tensor in the two-dimensional Ising spin model [7], we derive a complete expression including normalization that works under an arbitrary affine transformation. To our knowledge, our work is the first work that verifies the analytic formulas for the Ising CFT involving the EM tensor numerically with the spin variables. The details of the content will be given in a subsequent paper [8].
@@ 3 Partition functions
As is well known, the 2D Ising model has a fermionic description with the Majorana fermion [9, 10, 11]. For later discussion, we consider the Wilson-Majorana fermion action [12, 13] on the hexagonal lattice in a generalized setup with arbitrary couplings:
@@
The partition functions (4), (5) and (6) can be related to each other by loop expansions [14, 12, 13] even under various boundary conditions on the torus. For the discussion below, it suffices to quote the following identities:
@@ 4 Staggered lattice structure in the affine-transformed hexagonal lattice
In Ref. [13], it was shown that the continuum limit can be taken under nontrivial affine transformation with the couplings:
@@ 5 Lattice operators
We comment that both $\varepsilon(x)$ and $T_{k}(x)$ acquire divergent parts through loop diagrams (the divergent part of $\varepsilon(x)$ comes from the Wilson term). The divergent parts need to be subtracted when we consider disconnected components. This corresponds to taking the normal ordering. For the Ising CFT, its evaluation can be performed with the fermion system without statistical error [8]. In the numerical analysis below, all the lattice operators are regularized in this way.
@@ 6 Confirming the CFT formulas with Monte Carlo calculations
In this section, we calculate various expectation values that involve the EM tensor with the spin variable by performing Monte Carlo calculation. Ensembles are generated by combining the Wolff algorithm [12] with the local heat-bath algorithm. The details of the calculation will be described in Ref. [8]. We comment that that the signal of the EM tensor is noisy as the contributions from 1 (the diverging part), $\varepsilon(x)$, and $\partial\varepsilon(x)$ (the first descendant of $\varepsilon(x)$) need to be subtracted to single out the EM tensor components. The exact values are evaluated with the analytic expressions known in the Ising CFT [5, 6]. The modulus is set to a nontrivial value $\tau=1.2e^{4i\pi/9}$ below.
@@
We next consider the three-point function with the spin operators. This is the least noisy correlator involving the EM tensor. Recall that the conformal Ward identity for the primary fields $\phi_{i}(z,\bar{z})$ with the conformal weights $(h_{i},\tilde{h}_{i})$ is [5]:
@@ Acknowledgments
The authors thank Hidenori Fukaya, Okuto Morikawa, and Yusuke Namekawa for valuable discussions in LATTICE 2024. This work is partially supported by the Scientific Discovery through Advanced Computing (SciDAC) program, “Multiscale acceleration: Powering future discoveries in High Energy Physics” under FOA LAB-2580 funded by the U.S. DOE, Office of Science, and DOE under Award DE-SC0015845. The computation was performed on the Shared Computing Cluster (SCC) which is administered by Boston University’s Research Computing Services [16].